OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y − z, if x = 2, y = 3, and z = −6

OpenStudy (chaise):

4x − 2y − z, if x = 2, y = 3, and z = −6 Simple substitute in your value 4*2 - 2*3 - -6 = 8 - 6 - - 6 = 8 - 6 + 6 = 2 + 6 = 8 The answer is 8.